Suppose that A=(355)10 and B=(250)10 are 3-digit signed decimal numbers represented in 10’s complement.

(a) Perform the addition A+B in 10’s complement

(b) Does the addition in (a) result in an arithmetic overflow?
Answer #1

S=A+B = 355 + 250 = 605 in 10's complement form

S= 999-605 +1 = 395 in decimal form

b) no above addion doesnt have arithmetic overflow because there is no carry